WHICH NUMBER COMES NEXT IN THIS SERIES:5,7,11,19,35,?
Answer by stanbon(48568)

WHICH NUMBER COMES NEXT IN THIS SERIES:5,7,11,19,35,?
-----------------
7-5 = 2
11-7 = 4
19-11 = 8
35-19 = 16
? - 35 = 32
-------------------
? = 32+35 = 67
==========================
Cheers,
Stan H.